# Compressing Chain-of-Thought in LLMs via Step Entropy

### v1

@@ Appendix A Appendix
ACC (%) Tokens ACC (%) Tokens ACC (%) Tokens ACC (%) Tokens
DeepSeek-R1-7B $80.36$ 390 545390\,545 $88.17$ 1 851 9171\,851\,917 $63.33$ 475 303475\,303 $35.71$ 546 097546\,097
DeepSeek-R1-7B (Our) $80.82$ 388180 ($\downarrow$0.6%) $88.17$ 1302114 ($\downarrow$29.7%) $56.67$ 302784 ($\downarrow$36.3%) $35.71$ 344135 ($\downarrow$37.0%)
DeepSeek-R1-14B $82.64$ 374 109374\,109 $84.37$ 1 426 8651\,426\,865 $65.52$ 462 445462\,445 $58.62$ 540 003540\,003
DeepSeek-R1-14B (Our) $84.00$ 367060 ($\downarrow$1.9%) $82.16$ 990486 ($\downarrow$30.6%) $58.62$ 261167 ($\downarrow$43.5%) $51.72$ 325262 ($\downarrow$39.8%)
Qwen3-8B $94.46$ 4 027 7934\,027\,793 $91.37$ 3 569 2473\,569\,247 $79.31$ 628 097628\,097 $76.92$ 597 073597\,073
Qwen3-8B (Our) $94.39$ 3373301 ($\downarrow$16.2%) $91.13$ 2604619 ($\downarrow$27.0%) $81.48$ 346007 ($\downarrow$44.9%) $76.00$ 351499 ($\downarrow$41.1%)
QwQ-32B $94.09$ 2 610 1852\,610\,185 $92.35$ 2 977 6942\,977\,694 $79.13$ 637 318637\,318 $66.67$ 711 348711\,348
QwQ-32B (Our) $93.56$ 2148697 ($\downarrow$17.7%) $91.75$ 1989032 ($\downarrow$33.2%) $74.07$ 286324 ($\downarrow$55.1%) $65.52$ 403025 ($\downarrow$43.4%)
Table 5: Comparing the full COT baseline with our proposed step-entropy based pruning (Our) method, which prunes 80% of the lowest-entropy steps for DeepSeek-R1-7B, 14B, Qwen-8B and QwQ-32B. We conduct experiments to get the Pass@1 Accuracy(ACC)(%) and the number of Thinking Tokens (contains the Unicode characters) during the inference on GSM8k, Math500, AIME2024 and AIME2025.
@@
ACC (%) Tokens ACC (%) Tokens ACC (%) Tokens ACC (%) Tokens
QwQ-32B (Full COT) $94.09$ 2 610 1852\,610\,185 $92.35$ 2 977 6942\,977\,694 $79.13$ 637 318637\,318 $66.67$ 711 348711\,348
QwQ-32B (80%) $93.56$ 2148697 ($\downarrow$17.7%) $91.75$ 1989032 ($\downarrow$33.2%) $74.07$ 286324 ($\downarrow$55.1%) $65.52$ 403025 ($\downarrow$43.4%)
QwQ-32B (90%) $93.56$ 2010201 ($\downarrow$23.0%) $90.95$ 1505681 ($\downarrow$49.4%) $70.37$ 225711 ($\downarrow$64.6%) $66.67$ 193792 ($\downarrow$72.8%)
@@
To evaluate the domain generalizability of our step entropy-based compression method beyond mathematical reasoning, we conducted experiments on MMLU (Massive Multitask Language Understanding) benchmarks, specifically focusing on College Medicine and High School History tasks. Tables 8 and 7 present results for DeepSeek-R1-7B and QwQ-32B models respectively, comparing different compression levels (80%, 90% low-entropy step pruning) against full CoT and no-thinking baselines. The results reveal domain-specific compression characteristics that differ significantly from mathematical reasoning tasks, with both models showing varying degrees of compression tolerance across the two MMLU domains.
@@
DeepSeek-R1-7B (80%) $62.34$ 2127.9 ($\downarrow$18.6%) $64.32$ 1907.5 ($\downarrow$7.1%)
DeepSeek-R1-7B (90%) $62.34$ 2069.4 ($\downarrow$20.8%) $61.74$ 1936.2 ($\downarrow$5.7%)
@@
Broader Impact

### v2

@@ Appendix G Extended Experiments
Table 6: Comparing the full COT baseline with our proposed step-entropy based pruning (Our) method, which prunes 80% of the lowest-entropy steps for DeepSeek-R1-7B, 14B, Qwen-8B and QwQ-32B. We conduct experiments to get the Pass@1 Accuracy(ACC)(%) and the number of Average Thinking Tokens Per Answer (contains the Unicode characters) during the inference on GSM8k, Math500, AIME2024 and AIME2025.
@@
ACC (%) Avg Tokens ACC (%) Avg Tokens ACC (%) Avg Tokens ACC (%) Avg Tokens
DeepSeek-R1-7B 78.54 298.33 88.17 3703.83 63.33 15843.43 35.71 18203.23
DeepSeek-R1-7B (Our) 80.82 294.29 (↓1.3%) 88.17 2604.23 (↓29.7%) 56.67 10092.80 (↓36.3%) 35.71 11471.17 (↓37.0%)
DeepSeek-R1-14B 82.64 283.63 84.37 2853.73 65.52 15414.83 58.62 18000.10
DeepSeek-R1-14B (Our) 84.00 278.29 (↓1.9%) 82.16 1980.97 (↓30.6%) 58.62 8705.57 (↓43.5%) 51.72 10842.07 (↓39.8%)
Qwen3-8B 94.46 3053.67 91.37 7138.49 79.31 20936.57 76.92 19902.43
Qwen3-8B (Our) 94.39 2557.47 (↓16.2%) 91.13 5209.24 (↓27.0%) 81.48 11533.57 (↓44.9%) 76.00 11716.63 (↓41.1%)
QwQ-32B 94.09 1978.91 92.35 5955.39 79.13 21243.93 66.67 23711.60
QwQ-32B (Our) 93.56 1629.03 (↓17.7%) 91.75 3978.06 (↓33.2%) 74.07 9544.13 (↓55.1%) 65.52 13434.17 (↓43.4%)
@@
ACC (%) Avg Tokens ACC (%) Avg Tokens ACC (%) Avg Tokens ACC (%) Avg Tokens
QwQ-32B (Full COT) 94.09 1978.91 92.35 5955.39 79.13 21243.93 66.67 23711.60
QwQ-32B (80%) 93.56 1629.03 (↓17.7%) 91.75 3978.06 (↓33.2%) 74.07 9544.13 (↓55.1%) 65.52 13434.17 (↓43.4%)
QwQ-32B (90%) 93.56 1524.03 (↓23.0%) 90.95 3011.36 (↓49.4%) 70.37 7523.70 (↓64.6%) 66.67 6459.73 (↓72.8%)
@@
To evaluate the domain generalizability of our step entropy-based compression method beyond mathematical reasoning, we conducted experiments on MMLU (Massive Multitask Language Understanding) benchmarks, specifically focusing on College Medicine and High School History tasks. Tables 9 and 8 present results for DeepSeek-R1-7B and QwQ-32B models respectively, comparing different compression levels (80%, 90% low-entropy step pruning) against full CoT and no-thinking baselines. The results reveal domain-specific compression characteristics that differ significantly from mathematical reasoning tasks, with both models showing varying degrees of compression tolerance across the two MMLU domains.
@@
DeepSeek-R1-7B (80%) 62.34 2127.9 (↓18.6%) 64.32 1907.5 (↓7.1%)
DeepSeek-R1-7B (90%) 62.34 2069.4 (↓20.8%) 61.74 1936.2 (↓5.7%)
@@
Appendix H Extended Discussion on Limitations and Future Work
Our framework provides a robust method for CoT compression, but it also highlights several important areas for future investigation.
Sensitivity to Step Segmentation
[…]
Domain and Modality Generalizability
Our validation focuses primarily on mathematical reasoning and select knowledge-based tasks from MMLU. While we demonstrate strong performance in these areas, the transferability of step entropy as a universal metric for redundancy is not yet established for all reasoning types. The nature of redundancy in more open-ended, creative, or commonsense reasoning tasks may differ significantly. Therefore, an important direction for future work is to apply and adapt our framework to a broader range of domains. Furthermore, extending this method to multimodal reasoning, where steps may involve interpreting images or data visualizations, represents a challenging but valuable next frontier.
Adaptive and Model-Aware Compression
Our experiments establish an 80% low-entropy pruning threshold that works remarkably well across several models. However, we acknowledge that this fixed threshold may not be universally optimal. As shown in our own experiments, different models or problem domains can exhibit unique redundancy patterns. This suggests a need for more dynamic approaches. A promising area for future work is the development of an adaptive compression strategy. Such a system could dynamically adjust the pruning ratio $\kappa$ based on the model’s architecture, the specific problem’s complexity, or the reasoning domain, moving from a static threshold to a more intelligent, context-aware compression policy.
Appendix I Broader Impact